OpenStudy (anonymous):

How do I calculate this, I'm trying to explain my daughter's 4th grade homework to her. ..Adrian Dakota & Calvin made an average of 52 snacks Adrian & Dakota made an avereage of 61 snacks Dakota and Calvin made an average of 44 snacks. How many sancks did each individual make?

OpenStudy (anonymous):

we can write down 3 equations if that helps \[\frac{A+C+D}{3}=52\] \[\frac{A+D}{2}=61\] \[\frac{D+C}{2}=44\] and see if we can find what each of them are

OpenStudy (anonymous):

or maybe we can just think. since \[\frac{A+D}{2}=61\] that means \[A+D=121\] so \[\frac{A+C+D}{3}=\frac{C+121}{3}=52\] and so \[C+121 = 3\times 52=156\]therefore \[C=156-121=35\]

OpenStudy (anonymous):

now the others should be easy

OpenStudy (anonymous):

Thanks! I can't believe this is 4th grade math!!!

OpenStudy (anonymous):

yeah i was trying to think of a way to do this in our heads, but i cannot. there might be a simpler way without equations, but i can't think of it